OpenStudy (anonymous):

Can someone PLZ Describe the mathematic process of canceling like factors when working with rational expressions step-by-step IN WORDS.

OpenStudy (anonymous):

any example

OpenStudy (anonymous):

(3a^2-9a-12)/(6a^2+30a+24)

OpenStudy (anonymous):

3a^2 - 9a -12 = 3(a^2 -3a -4) = 3(a-4)(a+1) 6a^2 +30a +24 = 6(a^2 +5a +4) = 6(a+1)(a+4) \[\frac{3a^2 - 9a -12}{6a^2 +30a +24} = \frac{3(a+1)(a-4)}{6(a+1)(a+4)} = \frac{a-4}{2(a+4)}\]

jhonyy9 (jhonyy9):

3(a2 -3a-4)/6(a2 +5a +4)

OpenStudy (anonymous):

Can u Describe how u got that?

jhonyy9 (jhonyy9):

sheg this is your work

OpenStudy (anonymous):

i need it Describe step-by-step.

jhonyy9 (jhonyy9):

first steps are what i have wrote here little up that you can extracting 3 from all terms on numerator and 6 on denominator so is right ?

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok